# Enumerating Finite Braid Group Orbits on $SL_{2}(\mathbb{C})$-Character Varieties

### v2

@@ 3.1 Computing The Braid Group Orbit
The braid group can be presented as $B_{T+1}=\langle\sigma_{1},\dots,\sigma_{T}\mid\sigma_{i}\sigma_{i+1}\sigma_{i}=\sigma_{i+1}\sigma_{i}\sigma_{i+1},\sigma_{i}\sigma_{j}=\sigma_{j}\sigma_{i}\rangle$ where the first relation holds for $31\leq i\leq T-$ 3 and the second for $i-j\geq 2$. A generator $\sigma_{i}$ acts on $T+1$-tuples in the following way:
@@
Given a nice tuple $\mathbf{A}=[A_{1},A_{2},\dots,A_{T}]$, let $MC_{\lambda}(\mathbf{A})=[\tilde{A_{1}},\tilde{A_{2}},\dots,\tilde{A_{T}}]$ denote the middle convolution. If $(a_{1},a_{2},\dots,a_{T+1})\in(\mathbb{C}^{\times})^{T+1}$ is such that $a_{i}\tilde{A_{i}}\in SL_{2}(\mathbb{C})$ for $i=1,\dots T$ and $a_{T+1}((\tilde{A_{1}}\tilde{A_{2}}\cdots\tilde{A_{T}})^{-1})\in SL_{2}(\mathbb{C})$ with $\Pi^{T+1}a_{i}=1$, the tuple
$\mathbf{\hat{A}}=[a_{1}\tilde{A_{1}},a_{2}\tilde{A_{2}},\dots,a_{T+1}(\tilde{A_{1}}\tilde{A_{2}}\cdots\tilde{A_{T}})^{-1}]$
@@
When we say trace coordinate of a tuple $MC_{\lambda}(\mathbf{A})$ we will always mean the trace coordinates of the induced tuple $\mathbf{\hat{A}}$ unless otherwise specified.
We want to compute the orbit of induced tuples $\mathbf{\hat{A}}$ up to simultaneous conjugation. This is equivalent to computing the orbit of the point associated to $\mathbf{\hat{A}}$ in the character variety $\mathfrak{X}(\pi_{1}(\Sigma_{0,T+1}),SL_{2}(\mathbb{C}))$. Note that since all of our tuples come from FCRG and middle convolution is equivariant with the action of the braid group we already know all the orbits are finite.
@@
We first compute the signature of $\mathbf{M}$, and then apply each of the $T$ generators to the tuple $\mathbf{M}$. If $\sigma_{i}(\mathbf{M})$ has a new signature, then we store the tuple in $To\_Test$ and the signature in $Signatures$ and add $\mathbf{M}$ to $Tested$. We then apply $\sigma_{i}$ again to the (atmost $T$) new tuples in $To\_Test$ we got from the action on $\mathbf{M}$, once again storing only tuples which have new signatures. We repeat this process until the sets $To\_Test$ and $Tested$ are the same. The function returns the set $Tested$, which should contain only one representative per distinct signature, and can optionally return the set $Signatures$ as well. Note that here our definition of signature is the trace coordinates, which differs from the definition used in Tykhyy’s classification in terms of residues [31].
@@ 3.2 Finding Nice Tuples
Nice $T$-tuples correspond to MCG-finite $T+1$-tuples in the following way: given a nice tuple $[A_{1},\dots,A_{T}]$, we get a corresponding representation $[A_{1},\dots,A_{T},A_{T+1}]$ of the $T+1$-punctured sphere if we take $A_{T+1}=(A_{1}A_{2}\cdots A_{T})^{-1}$. Thus, it is enough to search each complex reflection group for nice tuples.
By Theorem 2.2.2, an eigenvalue of multiplicity $T-2$ for an element in $W$ must be a $d$-th root of unity for $d$ dividing at least $T-2$ degrees of $W$. Any reflection group can be constructed in Magma by providing the generating reflections; the irreducible finite complex reflection groups can be easily constructed in Magma by providing the Shephard-Todd number or the corresponding parameters for $G(m,p,r)$. In most cases Magma generates the group $W$ via the reflection representation over the trace field of $W$ [13, Table 1, Table 2].
@@
    Fix a representative $Q_{i}$ for each conjugacy class of reflections in $W$ and fix an ordering $[Q_{1},\cdots,Q_{n}]$.
@@
    Search $M$ for all sets such the matrices generate $W$ (means the tuple is irreducible) and the product of the matrices, for some order of multiplying, has an eigenvalue of multiplicity $T-2$. Each such set is then organized into a tuple based on the order of multiplication and stored in a list X of irreducible tuples.
@@
    Choose exemplars $\mathbf{A},\mathbf{B}$… for each type and compute the middle convolution $MC_{\lambda_{1}}(\mathbf{A}),MC_{\lambda_{2}}(\mathbf{A}),MC_{\lambda}(\mathbf{B})$… for suitable choice(s) of $\lambda$.
@@
    For each middle convolution of the exemplars, find a suitable character $[a_{1},\dots,a_{T+1}]$ and construct the induced tuple $\mathbf{\hat{A}}$
@@
    Compute the braid group orbit of $\mathbf{\hat{A}}$ and identify the $SL_{2}(\mathbb{C})$ subgroup generated by the matrices of $\mathbf{\hat{A}}$
@@
    Check the other exemplars and inverse tuple do not produce new orbits
@@
We recorded all the details for our exemplars and the middle convolution computations for each of the primitive reflection groups of rank 3 and 4. The middle convolution tables are listed in the Appendix. The types and exemplars are listed in the section for each group, as is the comparison tables with the other classifications. More details on what we use to compare with the existing classifications of [22] and [31] are given in Sections 5.1 and 5.2 .
@@ 4 Imprimitive Reflection Groups
Imprimitive reflection groups $G(m,p,n)$ are parameterized by $m,p,n\geq 1$ with $p\mid m$. These groups can be viewed as $n\times n$ generalized permutation matrices, which each row and column containing only one non-zero entry which is either 1 or an $\frac{m}{p}$-th root of unity. For $m>1$ and $(m,p,n)\neq(2,2,2)$ these groups are always irreducible, meaning they act irreducibly on $\mathbb{C}^{n}$ (see [6, Section 2]).
An element $A$ of $G(m,p,n)$ can be written as $A=[a_{1},a_{2},\dots,a_{n}\mid\sigma]$ for $\sigma\in S_{n}$ and $p\mid a_{1}+a_{2}\dots+a_{n}$. This element can be identified with the matrix where the $(i,\sigma(i))$-th entry of $A$ is $\zeta_{m}^{a_{i}}$ and all the other entries are 0 [29, Section 2.3]. Note that all these $a_{i}$ should be read modulo $m$ since we are working with powers of an $m$-th root of unity. Additionally, there is a natural surjective homomorphism $\varphi:G(m,p,n)\longrightarrow S_{n}$ given by sending $A\mapsto\sigma$.
@@
    Type 1: These are elements such that $\sigma=(ij)$ is a transposition with $a_{k}\equiv 0\mod m$ for $k\neq i,j$ and $a_{i}+a_{j}\equiv 0\mod m$. All reflections of Type 1 are of order 2 and lie in the subgroup $G(m,m,n)$ of $G(m,p,n)$. Using the notation from [29, Section 2.3] we write a Type 1 reflection as $s(i,j;a_{i})$ with the understanding that this is the same as $s(j,i;-a_{i})$.
@@
    When $p=1$, $S$ contains $n-1$ reflections of Type 1 and one reflection of Type 2 of order $m$
@@
Stated in [29, Lemma 3.2] with proofs in [27, Subsection 2.10 and Lemma 2.1] and [28, Lemma 2.2]. ∎
@@
This is true by the definition of reflections and the definition of the homomorphism. For a more thorough examination of related properties, we refer to [6, Proposition 2.2]. ∎
Given any set of reflections generating an irreducible imprimitive group the permutations associated to the reflections in the set must generate a transitive subgroup of $S_{n}$, or else the group cannot be irreducible. Note that in the case where the associated permutations are transpositions (i.e. the reflections are Type 1), the only transitive subgroup of $S_{n}$ generated by transpositions is $S_{n}$ itself [21, Lemma 2.13].
A subgroup of a reflection group generated by reflections is called a reflection subgroup. We modify some notation from [32] to state relevant theorems for identifying reflection subgroups in the imprimitive groups. Let $X=\{s(a_{h},a_{h+1};b_{k})\}$ be a set of reflections in $G(m,m,n)$. We can associate to $X$ the graph $\Gamma_{X}$ with nodes $\{a_{i}\}$ and edges $(a_{h},a_{h+1})$ for $s(a_{h},a_{h+1};b_{k})$ in $X$. Letting $\langle X\rangle$ denote the reflection group generated by the reflections in $X$, we see that this group is irreducible if and only if the graph $\Gamma_{X}$ is connected. Suppose $\Gamma_{X}$ is such that the graph contains exactly one cycle, given by $r$ edges corresponding to $\{s(a_{h},a_{h+1};b_{h})\}$ for $1\leq h\leq r$ for some $2\leq r\leq n$. Then we define the parameter $\delta(X)$ to be the absolute value of $\sum_{1}^{r}b_{h}$. If $p\neq m$, we say $X$ is a nice reflection set of $G(m,p,n)$ if $X$ contains exactly one reflection of Type 2.
@@
The first is proven in [27, Thm. 2.19] and the second is proven in [32, Theorem 3.10]. Note that in both these papers they refer to cycles as circles. The third statement is proven in [32, Theorem 3.9], where they define rooted trees to capture the presence of Type 2 reflections. We refer to these paper for more details on the graph theoretic interpretation of generating sets and their connection to reflection subgroups of the imprimitive reflection groups. ∎
@@
Among the imprimitive groups, we have to check for $n$-tuples and $n+1$-tuples in $G(m,p,n)$. By Theorem 4.0.1, the only possible groups that can contain nice $n$-tuples are the well-generated groups $G(m,1,n)$ and $G(m,m,n)$. We prove there are no nice $n$-tuples in either of these groups when $n\geq 5$ in Section 4.5. By results in [20, Lemma 5.3.2], we know that $n+1$-tuples can only exist in $G(m,p,n)$ for $n<5$. This work will fully analyze the existence of nice $n+1$-tuples for $52<n<$ 5.
@@ 4.1 Nice 3-Tuples in $G(m,p,3)$
In this case we are seeking nice tuples $[r_{1},r_{2},r_{3}]$ from $G(m,p,3)$ such that their product has a nontrivial eigenvalue of multiplicity 1. Note that since $\langle r_{1},r_{2},r_{3}\rangle=G(m,p,3)$ we know that $p=1$ or $p=m$ are the only possibilities as these are the only well-generated imprimitive groups (see Definition 2.2.1)
@@
In $G=G(m,1,3)$ for a nice tuple $[r_{1},r_{2},r_{3}]$ with $r_{1},r_{2}$ of Type 1 and $r_{3}$ of Type 2 with nontrivial eigenvalue $\zeta_{m}^{\prime}$ of order $m$, letting $A=r_{1}r_{2}r_{3}$, the eigenvalues of $A$ must be the distinct cube roots of $\zeta_{m}^{\prime}$.
@@
Thus we can compute the characteristic polynomial as $p_{A}(x)=-(x^{3}-\zeta_{m}^{a_{1}+a_{2}+a_{3}})$ and since the constant term of the characteristic polynomial is the determinant of $A$, we get that $\zeta_{m}^{\prime}=\zeta_{m}^{a_{1}+a_{2}+a_{3}}$. Thus $\lambda_{i}$ must be the distinct cube roots of $\zeta_{m}^{\prime}$. ∎
@@
In $G=G(m,m,3)$, if we have $r_{1}r_{2}r_{3}=A$ for reflections $r_{i}$ generating $G$, then the eigenvalues of $A$ are $\{\lambda^{-2},\lambda,-\lambda\}$
@@
$\zeta_{m}^{a_{k}}$ $=\lambda_{1}+\lambda_{2}+\lambda_{3}$
$-1$ $=\lambda_{1}\lambda_{2}\lambda_{3}$
The characteristic polynomial of $A$ is $p_{A}(x)=-(x^{2}-\zeta_{m}^{a_{i}+a_{j}})(x-\zeta_{m}^{a_{k}})$. Let $\lambda_{k}=\zeta_{m}^{a_{k}}$ so then we get that $\lambda_{i}=-\lambda_{j}$ since these are the roots of $x^{2}-\zeta_{m}^{a_{i}+a_{j}}$. Using the determinant equation we get $1=\lambda_{i}^{2}\lambda_{k}$ hence $\lambda_{i}^{2}=\lambda_{k}^{-1}=\zeta_{m}^{-a_{k}}=\zeta_{m}^{a_{i}+a_{j}}$ so both $\lambda_{i}$ and $\lambda_{j}$ square to the inverse of $\lambda_{k}$. ∎
@@
In this case we can consider reflections $[s(2,3;b_{1}),s(1,2;b_{2}),s(1,2;b_{3})]$ with $b_{2}$ and $b_{3}$ chosen so that $|b_{2}-b_{3}|$ is coprime to $m$. Taking $b_{1}=1,b_{2}=1$ and $b_{3}=0$, we get $\delta(X)=|b_{2}-b_{3}|$ is coprime to $m$ and so by Theorem 4.0.4, we know these reflections will generate $G(m,m,3)$. The characteristic polynomial of the product $r_{1}r_{2}r_{3}=A$ will in this case be $p_{A}(x)=-(x-\zeta_{m})(x^{2}-\zeta_{m}^{-1})$. Since the roots of the quadratic will be distinct, we know we will have at least two distinct eigenvalues, so the only worry is if the eigenvalues might be $\{\zeta_{m},\zeta_{m},1\}$ as then there is no nontrivial eigenvalue of multiplicity 1. If 1 is a root of the quadratic, then we know $\zeta_{m}^{-1}=1$ and $\zeta_{m}=1$ which is not possible. Hence, such nice tuples always exist. ∎

### v3

@@ 3.1 Computing The Braid Group Orbit
The braid group can be presented as $B_{T+1}=\langle\sigma_{1},\dots,\sigma_{T}\mid\sigma_{i}\sigma_{i+1}\sigma_{i}=\sigma_{i+1}\sigma_{i}\sigma_{i+1},\sigma_{i}\sigma_{j}=\sigma_{j}\sigma_{i}\rangle$ where the first relation holds for $1\leq i\leq T-3$ and the second for $i-j\geq 2$. A generator $\sigma_{i}$ acts on $T+1$-tuples in the following way:
@@
Given a nice tuple $\mathbf{A}=[A_{1},A_{2},\dots,A_{T}]$, let $MC_{\lambda}(\mathbf{A})=[\tilde{A_{1}},\tilde{A_{2}},\dots,\tilde{A_{T}}$] denote the middle convolution. If $(c_{1},c_{2},\dots,c_{T+1})\in(\mathbb{C}^{\times})^{T+1}$ is such that $c_{i}\tilde{A_{i}}\in SL_{2}(\mathbb{C})$ for $i=1,\dots T$ and $c_{T+1}((\tilde{A_{1}}\tilde{A_{2}}\cdots\tilde{A_{T}})^{-1})\in SL_{2}(\mathbb{C})$ with $\Pi^{T+1}c_{i}=1$, the tuple
$\mathbf{\hat{A}}=[c_{1}\tilde{A_{1}},c_{2}\tilde{A_{2}},\dots,c_{T+1}(\tilde{A_{1}}\tilde{A_{2}}\cdots\tilde{A_{T}})^{-1}]$
@@
Remark 3.1.3.
Note that over $\mathbb{C}$, it is always possible to find a scalar $c_{i}$ so that $\det(c_{i}A_{i})=1$ for any $A_{i}\in GL_{n}(\mathbb{C})$ as any $n$-th root of $\det A_{i}$ can be taken for $c_{i}$. As the matrices $\tilde{A_{1}},\dots,\tilde{A_{T+1}}$] multiply to the identity, there is always some choice of roots that multiply to the identity. In this work, we choose some scalars $(c_{1},\dots,c_{T+1})$ to construct our induced tuples. There may be multiple choices of such tuples of scalars, so our results should be considered up to choosing a different one.
When we say trace coordinate of a tuple $MC_{\lambda}(\mathbf{A})$ we will always mean the trace coordinates of the induced tuple $\mathbf{\hat{A}}$ unless otherwise specified. We want to compute the orbit of induced tuples $\mathbf{\hat{A}}$ up to simultaneous conjugation. This is equivalent to computing the orbit of the point associated to $\mathbf{\hat{A}}$ in the character variety $\mathfrak{X}(\pi_{1}(\Sigma_{T+1}),SL_{2}(\mathbb{C}))$. Note that since all of our tuples come from FCRG and middle convolution is equivariant with the action of the braid group we already know all the orbits are finite.
@@
We first compute the signature of $\mathbf{M}$, and then apply each of the $T$ generators to the tuple $\mathbf{M}$. If $\sigma_{i}(\mathbf{M})$ has a new signature, then we store the tuple in To_TestTo\_Test and the signature in $Signatures$ and add $\mathbf{M}$ to $Tested$. We then apply $\sigma_{i}$ again to the (atmost $T$) new tuples in To_TestTo\_Test we got from the action on $\mathbf{M}$, once again storing only tuples which have new signatures. We repeat this process until the sets To_TestTo\_Test and $Tested$ are the same. The function returns the set $Tested$, which should contain only one representative per distinct signature, and can optionally return the set $Signatures$ as well. Note that here our definition of signature is the trace coordinates (see Definition 2.3.2), which differs from the definition used in Tykhyy’s classification in terms of residues [Tyk].
@@ 3.2 Finding Nice Tuples
Nice $T$-tuples correspond to MCG-finite $T+1$-tuples in the following way: given a nice tuple $A_{1},\dots,A_{T}]$, we get a corresponding representation $A_{1},\dots,A_{T},A_{T+1}]$ of the $T+1$-punctured sphere if we take $A_{T+1}=(A_{1}A_{2}\cdots A_{T})^{-1}$. Thus, it is enough to search each complex reflection group for nice tuples.
By Theorem 2.2.2, an eigenvalue of multiplicity T2T-2 for an element in $W$ must be a $d$-th root of unity for $d$ dividing at least T2T-2 degrees of $W$. Any reflection group can be constructed in Magma by providing the generating reflections; the irreducible finite complex reflection groups can be easily constructed in Magma by providing the Shephard-Todd number or the corresponding parameters for $G(m,p,r)$. In most cases Magma generates the group $W$ via the reflection representation over the trace field of $W$ [feit2003, Table 1, Table 2].
@@
    Fix a representative $Q_{i}$ for each conjugacy class of reflections in $W$ and fix an ordering $Q_{1},\cdots,Q_{n}]$.
@@
    Search $M$ for all sets such that the matrices generate $W$ (means the tuple is irreducible) and the product of the matrices, for some order of multiplying, has an eigenvalue of multiplicity T2T-2. Each such set is then organized into a tuple based on the order of multiplication and stored in a list X of irreducible tuples.
@@
    Choose exemplars $\mathbf{A},\mathbf{B}$… for each type and compute the middle convolution
    $MC_{\lambda_{1}}(\mathbf{A}),MC_{\lambda_{2}}(\mathbf{A}),MC_{\lambda}(\mathbf{B})$… for suitable choice(s) of parameter $\lambda$.
@@
    For each middle convolution of the exemplars, find a suitable character $a_{1},\dots,a_{T+1}]$ and construct the induced tuple $\mathbf{\hat{A}}$.
@@
    Compute the braid group orbit of $\mathbf{\hat{A}}$ and identify the $SL_{2}(\mathbb{C})$ subgroup generated by the matrices of $\mathbf{\hat{A}}$.
@@
    Check the other exemplars and inverse tuple do not produce new orbits.
@@
We recorded all the details for our exemplars and the middle convolution computations for each of the primitive reflection groups of rank 3 and 4. The middle convolution tables are listed in the Appendix. The types and exemplars are listed in the section for each group, as is the comparison tables with the other classifications. More details on what we use to compare with the existing classifications of [T-L] and [Tyk] are given in Sections 5.1 and 5.2 .
@@ 4 Imprimitive Reflection Groups
Imprimitive reflection groups $G(m,p,n)$ are parameterized by $m,p,n\geq 1$ with $p\mid m$. These groups can be viewed as $n\times n$ generalized permutation matrices, which each row and column containing only one non-zero entry which is an $m$-th root of unity such that the product of all the nonzero entries is an $\frac{m}{p}$-th root of unity. For $m>1$ and $(m,p,n)\neq(2,2,2)$ these groups are always irreducible, meaning they act irreducibly on $\mathbb{C}^{n}$ (see [cohenfin, Section 2]).
An element $A$ of $G(m,p,n)$ can be written as $A=[a_{1},a_{2},\dots,a_{n}\mid\sigma]$ for $\sigma\in S_{n},a_{i}\in\mathbb{Z}$ and $p\mid a_{1}+a_{2}\dots+a_{n}$. This element can be identified with the matrix where the $(i,\sigma(i))$-th entry of $A$ is $\zeta_{m}^{a_{i}}$ and all the other entries are 0 [shiautos, Section 2.3]. Note that all these $a_{i}$ should be read modulo $m$ since we are working with powers of an $m$-th root of unity. Additionally, there is a natural surjective homomorphism $\varphi:G(m,p,n)\longrightarrow S_{n}$ given by sending $A\mapsto\sigma$.
@@
    Type 1: These are elements such that $\sigma=(ij)$ is a transposition with $a_{k}\equiv 0\mod m$ for $k\neq i,j$ and $a_{i}+a_{j}\equiv 0\mod m$. All reflections of Type 1 are of order 2 and lie in the subgroup $G(m,m,n)$ of $G(m,p,n)$. Using the notation from [shiautos, Section 2.3] we write a Type 1 reflection as $s(i,j;a_{i})$ with the understanding that this is the same as $s(j,i;-a_{i})$.
@@
    When $p=1$, $S$ contains n1n-1 reflections of Type 1 and one reflection of Type 2 of order $m$
@@
Stated in [shiautos, Lemma 3.2] with proofs in [shi2005mm, Subsection 2.10 and Lemma 2.1] and [shi2005mp, Lemma 2.2]. ∎
@@
This is true by the definition of reflections and the definition of the homomorphism. For a more thorough examination of related properties, we refer to [cohenfin, Proposition 2.2]. ∎
Given any set of reflections generating an irreducible imprimitive group the permutations associated to the reflections in the set must generate a transitive subgroup of $S_{n}$, or else the group cannot be irreducible. Note that in the case where the associated permutations are transpositions (i.e. the reflections are Type 1), the only transitive subgroup of $S_{n}$ generated by transpositions is $S_{n}$ itself [ld2009, Lemma 2.13].
A subgroup of a reflection group generated by reflections is called a reflection subgroup. We modify some notation from [wang2010] to state relevant theorems for identifying reflection subgroups in the imprimitive groups. Let $X=\{s(a_{h},a_{h+1};b_{h})\}$, $a_{h}\neq a_{h+1}$ be a subset of reflections in $G(m,m,n)$. We can associate to $X$ the graph $\Gamma_{X}$ with nodes $\{a_{i}\}$ and edges $(a_{h},a_{h+1})$ for $s(a_{h},a_{h+1};b_{h})$ in $X$, where we allow multi-edges but no self-loops (i.e no edges from a node to itself). Letting $\langle X\rangle$ denote the reflection group generated by the reflections in $X$, we see that this group is irreducible if and only if the graph $\Gamma_{X}$ is connected. Whenever $\Gamma_{X}$ is such that the graph contains exactly one cycle, we can define a parameter $\delta(X)$ as follows: given a cycle consisting of $2\leq r\leq n$ edges corresponding to reflections $\{s(a_{h},a_{h+1};b_{h})\}$ for $1\leq h\leq r$, take $\delta(X)$ to be the absolute value of $\sum_{1}^{r}b_{h}$.
If $p\neq m$, we say $X$ is a nice reflection set of $G(m,p,n)$ if $X$ contains exactly one reflection of Type 2. Note that in [wang2010] the presence of reflections of Type 2 are represented in the graph $\Gamma_{X}$ as “rooted” nodes, but as these do not affect the parameter $\delta(X)$, we do not need this terminology for our purposes.
@@
The first is proven in [shi2005mm, Thm. 2.19] and the second is proven in [wang2010, Theorem 3.10]. Note that in both these papers they refer to cycles as circles. The third statement is proven in [wang2010, Theorem 3.9], where they define rooted trees to capture the presence of Type 2 reflections. We refer to these paper for more details on the graph theoretic interpretation of generating sets and their connection to reflection subgroups of the imprimitive reflection groups. ∎
@@
Among the imprimitive groups, we have to check for $n$-tuples and $n+1$-tuples in $G(m,p,n)$. By Theorem 4.0.1, the only possible groups that can contain nice $n$-tuples are the well-generated groups $G(m,1,n)$ and $G(m,m,n)$. We prove there are no nice $n$-tuples in either of these groups when $n\geq 5$ in Section 4.5. By results in [LLL, Lemma 5.3.2], we know that $n+1$-tuples can only exist in $G(m,p,n)$ for $n<5$. This work will fully analyze the existence of nice $n+1$-tuples for $2<n<5$.
@@ 4.1 Nice 3-Tuples in $G(m,p,3)$
In this case we are seeking nice tuples $r_{1},r_{2},r_{3}]$ from $G(m,p,3)$ such that their product has a nontrivial eigenvalue of multiplicity 1. Note that since $\langle r_{1},r_{2},r_{3}\rangle=G(m,p,3)$ we know that $p=1$ or $p=m$ are the only possibilities as these are the only well-generated imprimitive groups (see Definition 2.2.1)
@@
In $G=G(m,1,3)$ for a nice tuple $r_{1},r_{2},r_{3}]$ with $r_{1},r_{2}$ of Type 1 and $r_{3}$ of Type 2 with nontrivial eigenvalue $\zeta_{m}^{\prime}$ of order $m$, letting $A=r_{1}r_{2}r_{3}$, the eigenvalues of $A$ must be the distinct cube roots of $\zeta_{m}^{\prime}$.
@@
Thus we can compute the characteristic polynomial as $char_{A}(x)=-(x^{3}-\zeta_{m}^{a_{1}+a_{2}+a_{3}})$ and since the constant term of the characteristic polynomial is the determinant of $A$, we get that $\zeta_{m}^{\prime}=\zeta_{m}^{a_{1}+a_{2}+a_{3}}$. Thus $\lambda_{i}$ must be the distinct cube roots of $\zeta_{m}^{\prime}$. ∎
@@
In $G=G(m,m,3)$, if we have $r_{1}r_{2}r_{3}=A$ for reflections $r_{i}$ generating $G$, then the eigenvalues of $A$ are $\{\lambda^{-2},\lambda,-\lambda\}$ for $\lambda$ a $2m$-th root of unity.
@@
$\lambda_{i}+\lambda_{j}=0$
$\lambda_{i}\lambda_{j}=-\zeta_{m}^{-a_{k}}$
This tells us that $\lambda_{i}=-\lambda_{j}$ and $\lambda_{i}^{2}=\zeta_{m}^{-a_{k}}$ so $\lambda_{i}$ must be a $2m$-th root of unity (not necessarily primitive). ∎
Remark 4.1.4.
Note that in the proof above and throughout this work, the convention for characteristic polynomial is $\det(A-xI)$.
@@
In this case we can consider reflections $s(2,3;b_{1}),s(1,2;b_{2}),s(1,2;b_{3})]$ with $b_{2}$ and $b_{3}$ chosen so that $|b_{2}-b_{3}|$ is coprime to $m$. Taking $b_{1}=1,b_{2}=1$ and $b_{3}=0$, we get $\delta(X)=|b_{2}-b_{3}|$ is coprime to $m$ and so by Theorem 4.0.4, we know these reflections will generate $G(m,m,3)$. The characteristic polynomial of the product $r_{1}r_{2}r_{3}=A$ will in this case be $char_{A}(x)=-(x-\zeta_{m})(x^{2}-\zeta_{m}^{-1})$. Since the roots of the quadratic will be distinct, we know we will have at least two distinct eigenvalues, so the only worry is if the eigenvalues might be $\{\zeta_{m},\zeta_{m},1\}$ as then there is no nontrivial eigenvalue of multiplicity 1. If 1 is a root of the quadratic, then we know $\zeta_{m}^{-1}=1$ and $\zeta_{m}=1$ which is not possible. Hence, such nice tuples always exist. ∎